OpenStudy (anonymous):

Can anyone help me with “Using the Rules of Exponents? If so, please show me the steps. I have two equations here is the first one division. Thank you. 27 ^-2/3 ÷ 27 ^-1/3

OpenStudy (jdoe0001):

$$ \cfrac{27^{-\frac{2}{3}}} {27^{-\frac{1}{3}}}\\ \color{red}{n^{-1}=\cfrac{1}{n}}\\ \cfrac{ \cfrac{1}{27^{\frac{2}{3}}} }{ \cfrac{1}{27^{\frac{1}{3}}} } \implies \cfrac{27^{\frac{1}{3}}} {27^{\frac{2}{3}}}\\ \color{red}{27 = 3^3}\\ \cfrac{(3^3)^{\frac{1}{3}}} {(3^3)^{\frac{2}{3}}}\implies \cfrac{(3)^{\frac{3}{3}}} {(3)^{\frac{6}{3}}}\implies \cfrac{3^1}{3^2} \implies \cfrac{1}{3} $$

OpenStudy (anonymous):

Hi Jdoe Thank you for your help. I got as far as 1/27. I am still trying to understand the exponents power rule. Thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Can you show me the steps to this equation? (a^1/2 b)^1/2 (ab ^1/2)

OpenStudy (jdoe0001):

$$ \cfrac{(a^{\frac{1}{2}}b)^{\frac{1}{2}}} {2ab^{\frac{1}{2}}} \large ? $$

OpenStudy (anonymous):

Can you show me the steps?

OpenStudy (jdoe0001):

well, first you multiply the exponential to the factors

OpenStudy (jdoe0001):

so \((a^{\frac{1}{2}}b)^{\frac{1}{2}} \) turns into \((a^{\frac{1}{2}}b)^{\frac{1}{2}} \implies a^{\frac{1}{2}\times \frac{1}{2}} b^{\frac{1}{2}}\)

OpenStudy (anonymous):

I worked on this yesterday with someone and was told ^1/2 x ^1/2 = 1^1/4

OpenStudy (jdoe0001):